# A Unified View of Group Fairness Tradeoffs Using Partial Information Decomposition

### v1

@@ Abstract
I Introduction
footnotetext: This paper has been published at the 2024 IEEE International Symposium on Information Theory (ISIT 2024), held in Athens, Greece.
The rapid infiltration of machine learning (ML) into high-stakes applications such as employment, education, finance, healthcare brings the promise of enhanced efficiency. However, this can also accompanied by escalating concerns about the disparate impact [1, 2, 3, 4, 5, 6] that these systems might cause on unprivileged groups based on sensitive attributes such as gender, race, age, nationality, etc. Several anti-discrimination legislations and ethical principles [1] are being actively put forth to ensure algorithmic fairness.
Existing literature has fostered a plethora of definitions, metrics, and scholarly debates about algorithmic fairness [7]. Central to the debate of quantifying fairness at a group level are three popular definitions, namely, statistical parity, equalized odds, and predictive parity [7, 8, 9]. Due to the multitude of fairness definitions available, it is often unclear which measure of fairness is most appropriate to adopt in a given setting [10]. Furthermore, it is also well-known that simultaneous satisfiability of these three fairness definitions is generally impossible [11, 12].
@@ I Introduction
Although previous studies have identified certain impossibilities among these fairness notions [11, 12], a detailed analysis focusing on the interrelationships among different measures of unfairness, specifically explaining when they will be in agreement and when they will be in disagreement leading to potential tradeoffs has received limited attention.
@@
Our research bridges this gap by leveraging Partial Information Decomposition (PID) [15], a body of work in information theory, to elucidate the exact relationship between different measures of unfairness. In particular, we consider information-theoretic quantifications [16] of the respective gaps from statistical parity, equalized odds, and predictive parity as our measures of unfairness. Using PID, we demonstrate the exact relationship between these three measures of unfairness in Proposition 1 (also see Fig. 3 for a pictorial illustration of the relationship between the measures of unfairness).
@@
PID enables us to provide a unified information-theoretic framework that is instrumental in establishing the fundamental limits and tradeoffs among these unfairness measures, particularly in the context of approximate fairness solutions when exact satisfiability of all three fairness definitions is not met. Furthermore, the impossibility among the three fairness definitions can also be derived from our result (see Theorem 1). We also identify and delineate the regions of agreement and disagreement among these three measures of unfairness (see Section III), providing insights on when there will be a tradeoff and when there will be no tradeoff among the measures of unfairness. We perform numerical simulations on the Adult dataset [17] to complement our theoretical results. Moreover, our work holds broader implications in fields such as algorithmic fairness auditing [18], where it can significantly contribute to the evaluation of fairness in ML models.
Related Works: Information-theoretic measures have been used to study group fairness in the fairness literature [19, 20, 16, 4, 5, 21, 22, 23, 24, 25, 26, 27]. Another related line of work is exploring trade-offs between fairness and accuracy [28, 29, 30, 31, 13, 32, 33, 14, 34, 35, 36, 37, 38, 39, 40].
@@
PID is recently gaining traction across various ML applications [41, 5, 4, 36, 42, 43, 44, 45, 46, 47]. It is particularly noteworthy in the realm of algorithmic fairness  [41, 5, 4, 36]. [5] leverages PID to dissect total disparity in decision-making into exempt and non-exempt components. Similarly, [36] employs PID to study the interplay between global and local fairness in federated learning. We also refer to [41] for a survey of PID in fairness and explainability. Understanding tradeoffs and agreement disagreement between unfairness measures using PID has not been studied. In this work, our objective is to develop a unified information-theoretic framework that effectively delineates the fundamental limits and trade-offs among the existing unfairness measures: statistical parity, equalized odds, and predictive parity.
II Preliminaries
@@ II-A Background on Partial Information Decomposition
We also note that defining any one of the PID terms suffices in obtaining the others. This is because of another relationship among the PID terms as follows [15]: $\mathrm{I}({Z;A})=\mathrm{Uni}({Z{:}A|B})+\mathrm{Red}({Z{:}A,B})$. Essentially $\mathrm{Red}({Z{:}A,B})$ is viewed as the sub-volume between $\mathrm{I}({Z;A})$ and $\mathrm{I}({Z;B})$ (see Fig. 1). Hence, $\mathrm{Red}({Z{:}A,B})=\mathrm{I}({Z;A})-\mathrm{Uni}({Z{:}A|B})$. Lastly, $\mathrm{Syn}({Z{:}A,B})=\mathrm{I}(Z;A,B)-\mathrm{Uni}({Z{:}A|B})-\mathrm{Uni}({Z{:}B|A})-\mathrm{Red}({Z{:}A,B})$ (can be obtained from (2) once both unique and redundant information has been defined).
The main results of our paper hold regardless of the specific definition of a given PID term. However, our experiments are based on the precise definition of $\mathrm{Uni}({Z{:}A|B})$ from [15].
@@
Operational meaning of Unique Information from Blackwell sufficiency: Unique information is closely tethered to Blackwell Sufficiency [48] in statistical decision theory. The concept of Blackwell sufficiency [48] from statistical decision theory helps characterize if a random variable $A$ is more informative than $B$ about $Z$ (also relates to stochastic degradation of channels [49, 50]). A channel $P_{B|Z}$ is Blackwell sufficient with respect to another channel $P_{A|Z}$ (also denoted as $B\geq_{Z}A$) if there exists a stochastic transformation $P_{A^{\prime}|B}$ such that the effective channel from $Z$ to $A^{\prime}$ is equivalent to the original channel from $Z$ to $A$ (see Fig. 2). The unique information $\mathrm{Uni}({Z{:}A|B})$ is 00 if and only if $P_{B|Z}$ is Blackwell sufficient with respect to $P_{A|Z}$ [15, 51, 50, 49]. Otherwise, $\mathrm{Uni}({Z{:}A|B})>0$, and it is viewed as a departure from Blackwell sufficiency, i.e., there exists a scenario where $A$ gives something unique about $Z$ that you can never get after degrading to $B$.
III Partial Information Decomposition of
the three Measures of Unfairness
@@ III Partial Information Decomposition of
We leverage PID to derive exact relationships among the three measures of unfairness. We decompose the statistical parity gap $I(Z;\hat{Y})$, equalized odds gap $I(Z;\hat{Y}|Y)$, and predictive parity gap $I(Z;Y|\hat{Y})$ into nonnegative overlapping terms. The significance of this decomposition is that it highlights regions where these measures are in agreement and disagreement. Fig. 3 provides a pictorial illustration of the overlaps between these three measures of unfairness.
@@
Proof Sketch: The proof relies on the nonnegativity of each of the PID terms (also recall Fig. 3). PID of $\mathrm{I}({Z;\hat{Y},Y})$ is expressed as $\mathrm{I}({Z;\hat{Y},Y})=\mathrm{Uni}({Z{:}\hat{Y}|Y})+\mathrm{Uni}({Z{:}Y|\hat{Y}})+\mathrm{Red}({Z{:}\hat{Y},Y})+\mathrm{Syn}({Z{:}\hat{Y},Y})$. Since each component in this decomposition is nonnegative, the presence of mutual information ($\mathrm{I}({Z;\hat{Y},Y})>0$) implies that at least one of these terms has a nonzero contribution. According to Proposition 1, each of these PID terms influences at least one unfairness measure. Therefore, the nonnegativity of any one of these terms results in at least one of the unfairness measures being nonzero. ∎
@@
TABLE I: Results of Regularizers on Different Measures of Unfairness
@@
IV Tradeoffs Between Unfairness Measures
We explore scenarios where models are trained with a focus on achieving any one specific fairness criterion and analyze its implications on the other two fairness notions. This applies to models that have been trained to achieve fairness either through in-processing techniques, such as adding fairness regularizers to the loss function, or through post-processing methods that adjust model outputs after training.
Theorem 3.
Furthermore, when $\mathrm{I}({Z;Y})=0$, it results in $\mathrm{Uni}({Z{:}Y|\hat{Y}})+\mathrm{Red}({Z{:}\hat{Y},Y})=0$, leading to each of those individual terms being zero, i.e., $\mathrm{Uni}({Z{:}Y|\hat{Y}})=0$ and $\mathrm{Red}({Z{:}\hat{Y},Y})=0$. Therefore, $\mathrm{I}({Z;Y|\hat{Y}})=\mathrm{Syn}({Z{:}\hat{Y},Y})\ =\mathrm{I}({Z;\hat{Y}|Y})$. ∎
Theorem 4.
Theorem 5.
Proof Sketch Given that Equalized Odds is met, we have $\mathrm{I}({Z;\hat{Y}|Y})=\mathrm{Uni}({Z{:}\hat{Y}|Y})+\mathrm{Syn}({Z{:}\hat{Y},Y})=0$. Consequently, from nonnegativity, both the terms $\mathrm{Uni}({Z{:}\hat{Y}|Y})$ and $\mathrm{Syn}({Z{:}\hat{Y},Y})$ are 0. Statistical Parity gap simplifies to $\mathrm{I}({Z;\hat{Y}})=\mathrm{Uni}({Z{:}\hat{Y}|Y})+\mathrm{Red}({Z{:}\hat{Y},Y})=\mathrm{Red}({Z{:}\hat{Y},Y})$, and the Predictive Parity gap is expressed as $\mathrm{I}({Z;Y|\hat{Y}})=\mathrm{Uni}({Z{:}Y|\hat{Y}})+\mathrm{Syn}({Z{:}\hat{Y},Y})=\mathrm{Uni}({Z{:}Y|\hat{Y}})$. Hence, $\mathrm{I}({Z;Y})=\mathrm{Uni}({Z{:}Y|\hat{Y}})+\mathrm{Red}({Z{:}\hat{Y},Y})=\mathrm{I}({Z;\hat{Y}})+\mathrm{I}({Z;Y|\hat{Y}})$. Since, $\mathrm{I}({Z;Y})$ is fixed for a dataset, an increase in the statistical parity gap leads to a decrease in the predictive parity gap, and vice versa. ∎
@@ IV Tradeoffs Between Unfairness Measures
Experimental Demonstrations. In this section, we provide an experimental demonstration on the Adult dataset [17] to validate our theoretical findings. The classification task for this dataset involves predicting whether an individual’s income exceeds $50$K per year, using features such as occupation, education, etc. We use gender as a sensitive attribute.
@@
We train a neural network consisting of a sequence of layers: the input layer is followed by three hidden layers, each with $32$ units and ReLU activation, and concludes with a single output layer using a sigmoid activation function. Training is conducted using a batch size of $512$, and the Adam optimizer with a learning rate of $0.01$. We apply various fairness regularizers and measure the unfairness as well as their decomposition (results are summarized in Table.I). We use the dit package [53] for PID computation and FairTorch [54] for fairness regularizer implementation.
@@
A key observation in our analysis is that $\mathrm{I}({Z;Y})$ consistently measures 0.037 using the Adult dataset. This mass does not decrease across various models since it only depends on the dataset. The PID terms in $\mathrm{I}({Z;Y})$, i.e., $\mathrm{Uni}({Z{:}Y|\hat{Y}})$ and $\mathrm{Red}({Z{:}\hat{Y},Y})$ contribute to either predictive parity or statistical parity gap. Also, when statistical parity is achieved (scenario with SP regularizer), the predictive parity gap is greater than the equalized odds gap. Also due to the impossibility of attaining zero unfairness with all the measures (see scenario with SP, EO, and PP regularizers), the mass typically moves to $\mathrm{Uni}({Z{:}Y|\hat{Y}})$, contributing to the predictive parity.

### v2

@@ Abstract
1 Introduction
footnotetext: Presented at the IEEE International Symposium on Information Theory (ISIT 2024) in Athens, Greece.footnotetext: The authors are with the Department of Electrical and Computer Engineering, University of Maryland College Park, United States 20742. Author contacts: fhamman@umd.edu, sanghamd@umd.edufootnotetext: This work was supported in part by NSF CAREER Award 2340006.
The increasing adoption of machine learning (ML) in high-stakes applications such as employment, finance, healthcare, etc, promises enhanced efficiency and improved decision-making. However, this widespread reliance on ML systems has escalated concerns about the disparate impact [1, 2, 3, 4, 5, 6, 7, 8, 9] that these systems might cause on unprivileged groups based on sensitive attributes such as gender, race, age, etc. Several anti-discrimination laws and ethical principles [1] are being actively put forth to ensure algorithmic fairness.
Existing literature has studied a plethora of definitions, metrics, and scholarly debates about algorithmic fairness [2, 3]. Central to the debate of quantifying fairness at a group level are three popular definitions, namely, statistical parity, equalized odds, and predictive parity [2, 10, 3]. Due to the multitude of fairness definitions available, it is often unclear which measure of fairness is most appropriate to adopt in a given setting [11]. Furthermore, it is also well-known that simultaneous satisfiability of these three fairness definitions is generally impossible [12, 13, 14].
@@ 1 Introduction
Although previous studies have identified certain impossibilities among these fairness notions [13, 12, 14], a detailed analysis focusing on the interrelationships among different measures of unfairness, specifically explaining when they will be in agreement and when they will be in disagreement leading to potential tradeoffs has received limited attention.
@@
Our research bridges this gap by leveraging Partial Information Decomposition (PID) [17], a body of work in information theory, to elucidate the exact relationship between different measures of unfairness. In particular, we consider information-theoretic quantifications [18] of the respective gaps from statistical parity, equalized odds, and predictive parity as our measures of unfairness. Using PID, we demonstrate the exact relationship between these three measures of unfairness in Proposition 1. We also refer to Fig. 1 and Fig. 4 for a pictorial illustration of the relationship between the measures of unfairness.
@@
PID enables us to provide a unified information-theoretic framework that is instrumental in establishing the fundamental limits and tradeoffs among these unfairness measures, particularly in the context of approximate fairness solutions when exact satisfiability of all three fairness definitions is not met. Furthermore, the impossibility among the three fairness definitions can also be derived from our result (see Theorem 1). We also identify and delineate the regions of agreement and disagreement among these three measures of unfairness (see Section 3), providing insights on when there will be a tradeoff and when there will be no tradeoff among the measures of unfairness. We perform numerical simulations on the Adult dataset [19] to complement our theoretical results. Moreover, our work holds broader implications in fields such as algorithmic fairness auditing [8], where it can significantly contribute to the evaluation of fairness in ML models.
Figure 1: Illustrates the decomposition of mutual information $\mathrm{I}({Z;\hat{Y},Y})$ using the chain rule. (left) shows the decomposition into Statistical Parity and Predictive Parity. (right) shows the decomposition into $\mathrm{I}({Z;Y})$ and Equalized Odds. No further insights into the overlapping regions of these measures, highlighting the need for measures to capture the nuanced interactions between fairness measures.
1.1 Related Works
Existing literature has introduced several foundational definitions and methods for fairness in machine learning (see some comprehensive overviews and surveys [4, 5, 3, 2]). Another line of work focuses on exploring trade-offs between fairness and accuracy [20, 15, 21, 22, 16, 23, 24, 25, 26, 27, 28, 29, 30, 31].
Early works on impossibility results among different fairness measures highlight the challenges of simultaneously satisfying multiple group fairness criteria [13, 12, 32, 33, 14]. Notably, [34] challenges the practical implications of the impossibility theorem, showing that fairness across multiple criteria is more achievable than previously believed. Similarly, [31] presents an integer-programming-based framework for optimizing post-processing methods to simultaneously satisfy multiple fairness criteria under small violations while maintaining a minimal reduction in model performance. However, the nuanced interrelationships among different measures of unfairness, specifically explaining when they will be in agreement and when they will be in disagreement leading to potential tradeoffs has received limited attention.
Information-theoretic measures have been used to quantify group fairness in existing literature [35, 36, 18, 7, 37, 38, 39, 40, 41, 42, 43, 44]. For instance, [18, 7] have already used mutual information and conditional mutual information to quantify statistical parity, equalized odds, and predictive parity. Closely related to our work, [30] explores whether fairness measures can be gradually compatible using these information-theoretic quantifications and chain rule, also showing that some fairness criteria can be simultaneously improved through fairness-regularized predictors. Our work dives deeper into this nuanced tradeoff among the three group fairness measures leveraging a body of work in information theory called Partial Information Decomposition (PID) that goes beyond chain rule, delineating regions of agreement and disagreement.
@@ 1.1 Related Works
PID is recently gaining traction across various ML applications [45, 37, 7, 25, 46, 47, 48, 49, 50, 51]. It is particularly noteworthy in the realm of algorithmic fairness [45, 37, 7, 25]. Prior work [37, 7] leverages PID to dissect total disparity in decision-making into exempt and non-exempt components depending on which features they came from. Another work [25] leverages PID to analyze the trade-offs between global and local fairness in a federated learning setting, identifying three sources of unfairness, and formulates a convex optimization problem to define the theoretical limits of accuracy and fairness trade-offs. We also refer to [45] for a survey of PID in fairness and explainability. Understanding tradeoffs and agreement disagreement between the three canonical group fairness measures, namely statistical parity, equalized odds, and predictive parity, using PID has not been studied. We aim to develop a unified information-theoretic framework that effectively formalizes the fundamental limits and trade-offs among these three unfairness measures.
2 Preliminaries
@@ 2.1 Background on Partial Information Decomposition
We also note that defining any one of the PID terms suffices in obtaining the others. This is because of another relationship among the PID terms as follows [17]: $\mathrm{I}({Z;A})=\mathrm{Uni}({Z{:}A|B})+\mathrm{Red}({Z{:}A,B})$. Essentially $\mathrm{Red}({Z{:}A,B})$ is viewed as the sub-volume between $\mathrm{I}({Z;A})$ and $\mathrm{I}({Z;B})$ (see Fig. 2). Hence, $\mathrm{Red}({Z{:}A,B})=\mathrm{I}({Z;A})-\mathrm{Uni}({Z{:}A|B})$. Lastly, $\mathrm{Syn}({Z{:}A,B})=\mathrm{I}(Z;A,B)-\mathrm{Uni}({Z{:}A|B})-\mathrm{Uni}({Z{:}B|A})-\mathrm{Red}({Z{:}A,B})$ (can be obtained from (2) once both unique and redundant information has been defined). The main results of our paper hold regardless of the specific definition of a given PID term. However, our experiments are based on the precise definition of $\mathrm{Uni}({Z{:}A|B})$ from [17].
@@
Operational meaning of Unique Information from Blackwell sufficiency: Unique information is closely tethered to Blackwell Sufficiency [52] in statistical decision theory. The concept of Blackwell sufficiency [52] from statistical decision theory helps characterize if a random variable $A$ is more informative than $B$ about $Z$ (also relates to stochastic degradation of channels [53, 51]). A channel $P_{B|Z}$ is Blackwell sufficient with respect to another channel $P_{A|Z}$ (also denoted as $B\geq_{Z}A$) if there exists a stochastic transformation $P_{A^{\prime}|B}$ such that the effective channel from $Z$ to $A^{\prime}$ is equivalent to the original channel from $Z$ to $A$ (see Fig. 3). The unique information $\mathrm{Uni}({Z{:}A|B})$ is 00 if and only if $P_{B|Z}$ is Blackwell sufficient with respect to $P_{A|Z}$ [17, 51, 53]. Otherwise, $\mathrm{Uni}({Z{:}A|B})>0$, and it is viewed as a departure from Blackwell sufficiency, i.e., there exists a scenario where $A$ gives something unique about $Z$ that you can never get after degrading to $B$.
@@
3 Decomposition of the Measures of Unfairness
@@
We leverage PID to derive exact relationships among the three measures of unfairness. We decompose the statistical parity gap $I(Z;\hat{Y})$, equalized odds gap $I(Z;\hat{Y}|Y)$, and predictive parity gap $I(Z;Y|\hat{Y})$ into nonnegative overlapping terms. The significance of this decomposition is that it highlights regions where these measures are in agreement and disagreement. Fig. 1 and Fig. 4 provides a pictorial illustration of the overlaps between these three measures of unfairness.
@@
Proof Sketch: The proof relies on the nonnegativity of each of the PID terms (also recall Fig. 4). PID of $\mathrm{I}({Z;\hat{Y},Y})$ is expressed as $\mathrm{I}({Z;\hat{Y},Y})=\mathrm{Uni}({Z{:}\hat{Y}|Y})+\mathrm{Uni}({Z{:}Y|\hat{Y}})+\mathrm{Red}({Z{:}\hat{Y},Y})+\mathrm{Syn}({Z{:}\hat{Y},Y})$. Since each component in this decomposition is nonnegative, the presence of mutual information ($\mathrm{I}({Z;\hat{Y},Y})>0$) implies that at least one of these terms is nonzero. According to Proposition 1, each of these PID terms influences at least one unfairness measure. Therefore, the nonnegativity of any one of these terms results in at least one of the unfairness measures being nonzero. ∎
@@
Figure 5: (left) Illustrates Theorem 3, showing that when Statistical Parity is satisfied, the Predictive Parity gap is greater than or equal to the Equalized Odds gap, and if $\mathrm{I}({Z;Y}){=}0$, then $\mathrm{I}({Z;Y|\hat{Y}}){=}\mathrm{I}({Z;\hat{Y}|Y})$. (right) visualizes Theorem 5 illustrating that when Equalized Odds is satisfied and $\mathrm{I}({Z;Y}){>}0$, there is an inverse relationship (tradeoff) between Statistical Parity and Predictive Parity ($\mathrm{I}({Z;\hat{Y}}){=}\mathrm{I}({Z;Y}){-}\mathrm{I}({Z;Y|\hat{Y}})$) since $\mathrm{I}({Z;Y})$ is fixed.
4 Tradeoffs Between Unfairness Measures
First, we will explore scenarios where models are trained with a focus on achieving any one specific fairness criterion and analyze its implications on the other two fairness notions. This applies to models that have been trained to achieve fairness either through in-processing techniques, such as adding fairness regularizers to the loss function, or through post-processing methods that adjust model outputs after training.
Theorem 3.
Furthermore, when $\mathrm{I}({Z;Y})=0$, it results in $\mathrm{Uni}({Z{:}Y|\hat{Y}})+\mathrm{Red}({Z{:}\hat{Y},Y})=0$, leading to each of those individual terms being zero, i.e., $\mathrm{Uni}({Z{:}Y|\hat{Y}})=0$ and $\mathrm{Red}({Z{:}\hat{Y},Y})=0$. Therefore, $\mathrm{I}({Z;Y|\hat{Y}})=\mathrm{Syn}({Z{:}\hat{Y},Y})\ =\mathrm{I}({Z;\hat{Y}|Y})$ (see Fig. 5 for an illustration). ∎
Theorem 4.
Theorem 5.
Proof Sketch Given that Equalized Odds is met, we have $\mathrm{I}({Z;\hat{Y}|Y})=\mathrm{Uni}({Z{:}\hat{Y}|Y})+\mathrm{Syn}({Z{:}\hat{Y},Y})=0$. Consequently, from nonnegativity, both the terms $\mathrm{Uni}({Z{:}\hat{Y}|Y})$ and $\mathrm{Syn}({Z{:}\hat{Y},Y})$ are 0. Statistical Parity gap simplifies to $\mathrm{I}({Z;\hat{Y}})=\mathrm{Uni}({Z{:}\hat{Y}|Y})+\mathrm{Red}({Z{:}\hat{Y},Y})=\mathrm{Red}({Z{:}\hat{Y},Y})$, and the Predictive Parity gap is expressed as $\mathrm{I}({Z;Y|\hat{Y}})=\mathrm{Uni}({Z{:}Y|\hat{Y}})+\mathrm{Syn}({Z{:}\hat{Y},Y})=\mathrm{Uni}({Z{:}Y|\hat{Y}})$. Hence, $\mathrm{I}({Z;Y})=\mathrm{Uni}({Z{:}Y|\hat{Y}})+\mathrm{Red}({Z{:}\hat{Y},Y})=\mathrm{I}({Z;\hat{Y}})+\mathrm{I}({Z;Y|\hat{Y}})$. Since, $\mathrm{I}({Z;Y})$ is fixed for a dataset, an increase in the statistical parity gap leads to a decrease in the predictive parity gap, and vice versa (see Fig. 5 for an illustration). ∎
5 Experimental Demonstrations
Table 1: Results of Regularizers on Different Measures of Unfairness
@@ 5 Experimental Demonstrations
In this section, we provide an experimental demonstration on the Adult dataset [19] to validate our theoretical findings. The classification task for this dataset involves predicting whether an individual’s income exceeds $50$K per year, using features such as occupation, education, etc. We use gender as a sensitive attribute.
@@
We train a neural network consisting of a sequence of layers: the input layer is followed by three hidden layers, each with $32$ units and ReLU activation, and concludes with a single output layer using a sigmoid activation function. Training is conducted using a batch size of $512$, and the Adam optimizer with a learning rate of $0.01$. We apply various fairness regularizers and measure the unfairness as well as their decomposition (results are summarized in Table.1). We use the dit package [55] for PID computation and FairTorch [56] for fairness regularizer implementation.
@@
A key observation in our analysis is that $\mathrm{I}({Z;Y})$ consistently measures 0.037 using the Adult dataset. This mass does not decrease across various models since it only depends on the dataset. The PID terms in $\mathrm{I}({Z;Y})$, i.e., $\mathrm{Uni}({Z{:}Y|\hat{Y}})$ and $\mathrm{Red}({Z{:}\hat{Y},Y})$ contribute to either predictive parity or statistical parity gap. When statistical parity is achieved (scenario with SP regularizer), the predictive parity gap is greater than the equalized odds gap. Also due to the impossibility of attaining zero unfairness with all the measures (see scenario with SP, EO, and PP regularizers), the mass typically moves to $\mathrm{Uni}({Z{:}Y|\hat{Y}})$, contributing to the predictive parity. The experiments confirm the theoretical insights by illustrating the inherent trade-offs between fairness measures. Specifically, even with fairness regularizers, achieving zero unfairness across all metrics is impossible due to the fixed information content in the dataset, i.e., $\mathrm{I}({Y;Z})$. The movement of unfairness mass between statistical parity and predictive parity highlights the difficulty of balancing fairness across multiple criteria, reinforcing the necessity of carefully considering these trade-offs in real-world applications.
6 Conclusion